### v1

@@ 2.5 DFT validations
We conduct a DFT validation of DeepRelax’s performance by comparing DFT calculated energy with the energy predicted by our model. To assess the predictive capability of our model under challenging conditions, we remove unrelaxed structures from the X-Mn-O test set which are structurally similar to their corresponding final counterparts using Pymatgen’s Structure_matcher function. From the remaining test set ($N=1\text{,}007$), we randomly selected 100 samples. Figure 4 (a) depicts the deviation distribution for the selected unrelaxed structures, which closely aligns with that of the complete test set, thereby affirming the representativeness of the selected subset.
Figure 4: (a) Histograms illustrating the density of deviations from the relaxed structure for the entire sample set and a selected subset of 100 samples, which demonstrates the subset’s representativeness. (b) Box plots illustrating the energy distributions for unrelaxed, DFT-relaxed, and DeepRelax-predicted structures, with the latter two exhibiting similar medians and spreads, confirming the accuracy of DeepRelax in approximating energetically favorable structures.
Subsequently, we employ DeepRelax to predict the relaxed structures for these samples. The energies of the unrelaxed, relaxed, and DeepRelax predicted structures are then calculated using DFT, as detailed in Subsection 4.9. Figure 4 (b) shows box plots of the energy distributions for the three structure categories. The predicted and relaxed structures have negative energy, and exhibit similar medians and interquartile ranges, suggesting that DeepRelax accurately predicts structures nearing their final ground states. This close alignment validates the model’s accuracy in predicting energetically favorable structures. The MAE in energy is significantly reduced by an order of magnitude from 32.51 for unrelaxed versus relaxed structures to only 5.97 for predicted versus relaxed structures.
@@ 2.6 Analysis of uncertainty
A critical challenge in integrating artificial intelligence into material discovery is establishing human trust in AI models. Current deep learning models typically offer reliable predictions only within the chemical space covered by their training datasets, known as the applicability domainyu2022uncertainty . Predictions for samples outside this domain can be unreliable. Thus, uncertainty quantification has become vital, enabling more trustworthy AI applications by quantifying prediction confidence levels, thereby aiding researchers in decision-making and experimental planning.
Figure 5: Uncertainty quantification. Hexagonal binning plots comparing global uncertainty with distance MAE for the (a) X-Mn-O, (b) MP, and (c) C2DB datasets. Parts (d) and (e) of the figure illustrate the local uncertainty in the bond lengths of the predicted structures for $\rm Sr_{2}Mn_{2}O_{4}$ and $\rm Mg_{1}Mn_{1}O_{3}$, respectively, demonstrating the correlation between bond length prediction errors and their associated uncertainties.
To validate the efficacy of our proposed uncertainty quantification in reflecting the confidence level of model predictions, we compute Spearman’s rank correlation coefficient between the predicted distance error and its associated global uncertainty. Figures 5 (a), (b), and (c) show the hexagonal binning plots of global uncertainty against distance MAE for the X-Mn-O, MP, and C2DB datasets, respectively. Correlation coefficients of 0.95, 0.83, and 0.88 for these datasets demonstrate a strong correlation between predicted error and predicted global uncertainty. Figures 5 (d) and (e) present the local (bond-level) uncertainty visualization for two predicted structures, illustrating the correlation between predicted bond length error and associated uncertainty. These results indicate that the model’s predicted uncertainty is a reliable indicator of the predicted structure’s accuracy.
Table 4: Ablation study results showcasing the impact of PCOE and local uncertainty estimation on model performance.
@@
DeepRelaxP 0.121 0.147 0.063 3.563
DeepRelaxU 0.142 0.171 0.064 3.539
@@ 2.7 Ablation study
DeepRelax’s technical contributions are twofold: it utilizes periodic cell offset encoding (PCOE) for handling PBCs explicitly, and it employs a method for estimating local uncertainty (i.e., bond-level uncertainty).
@@
  • Vanilla: Excludes both PCOE and local uncertainty estimation.
  • DeepRelaxP: Integrates PCOE but omits local uncertainty estimation.
  • DeepRelaxU: Implements local uncertainty estimation but not PCOE.
Table 4 demonstrates that DeepRelaxP attains a significant performance enhancement over the Vanilla model, suggesting the PCOE contributes greatly to model performance. On the other hand, DeepRelaxU shows a more modest improvement, which indicates the added value of local uncertainty estimation. Overall, DeepRelax shows a 25.16% improvement in coordinate MAE and a 20.00% advancement in bond length MAE over the Vanilla model. These comparative results underscore the combined effectiveness of positional PCOE and local uncertainty estimation in our final DeepRelax model.
@@ 3 Discussion
The rapid advancement of generative models like CDVAExie2021crystal , PGCGMzhao2023physics , and MatterGenzeni2023mattergen , has opened avenues for the prolific generation of hypothetical materials with potentially desirable properties, such as 2.2 million new materials recently discovered by Google DeepMind. Clearly, it is impossible to relax such a huge number of structures using the traditional ab initio method, and it is also very difficult using the iterative ML relax models. For example, we further compare the efficiency between DeepRelax and M3GNet, a leading iterative ML model. DeepRelax offers a substantial speed advantage, being approximately 100 times faster than M3GNet (see SM Section 3, Table S1). Based on this estimation, to relax the 2.2 million new materials, our DeepRelax model only needs around 100 hours or 4 days, while M3GNet will take around 400 days. Moreover, our DeepRelax model supports parallel GPU processing, which can further significantly reduce computer time. Overall, we introduce a fast, scalable, and reliable deep generative model, DeepRelax, for direct structural relaxation. Despite its advancements, opportunities for further improvement remain, which we explore in subsequent discussions.
@@
Figure 6: The architecture of PaEGNN. (a) Illustrates the multi-graph representation designed to capture atomic interactions across cell boundaries in periodic structures. (b) Depicts an overview of PaEGNN, comprising four layers, each consisting of message passing and message updating phases. The specifics of these phases are further detailed in (d) and (e), respectively.
@@ 4 Methods
4.1 Periodic structure of materials
Any material structure can be conceptualized as a periodic arrangement of atoms in 3D spacexie2021crystal . This periodicity is typically captured by identifying a repeating unit, known as a unit cell, that effectively represents the material’s infinite structure. Such a unit cell, containing $N$ atoms, can be fully characterized by three components:
@@ 4.1 Periodic structure of materials
where $(k_{1},k_{2},k_{3})$ are the cell offsets used to replicate the unit cell across the 3D space.
4.2 Multi-graph representation for periodic structure
Multi-graphs offer an intuitive way to represent crystal structures under periodic boundary conditions (PBCs)xie2021crystal , as depicted in Figure 6 (a). These graphs can be effectively processed by graph neural networks (GNNs) through graph convolutions or message passing, which simulate many-body interactionsgasteiger2021gemnet ; li2022deep ; zhong2023transferable ; xie2018crystal ; park2020developing ; schutt2017schnet ; chen2019graph ; gasteiger_dimenet_2020 ; choudhary2021atomistic ; chen2022universal ; deng2023chgnet ; unke2021spookynet ; pablo2023fast ; batzner20223 ; banik2023cegann ; unke2019physnet ; wang2022comenet ; zhang2022interpretable ; gilmer2017neural ; louis2020graph ; fuchs2020se ; brandstetter2021geometric ; satorras2021n ; dai2021graph ; omee2022scalable ; schutt2021equivariant ; haghighatlari2022newtonnet . Formally, we define a multi-graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ to encode these periodic structures. Here, $\mathcal{V}=\{v_{1},...,v_{N}\}$ represents the set of nodes (atoms), and =$\mathcal{E}=\{e_{ij,(k_{1},k_{2},k_{3})}|i,j\in\{1,...,N\},k_{1},k_{2},k_{3}\in\mathbb{Z}\}$ signifies the set of edges (bonds). The edge $e_{ij,(k_{1},k_{2},k_{3})}$ denotes a directed connection from node $i$ in the original unit cell to node $j$ in the cell translated by $k_{1}\vec{\bm{l}}_{1}+k_{2}\vec{\bm{l}}_{2}+k_{3}\vec{\bm{l}}_{3}$. Nodes are interconnected with their nearest neighbors within a cutoff distance $D$ (set to $6\ \rm\AA$ in our study).
@@ 4.3 Periodicity-aware equivariant graph neural network
PaEGNN iteratively updates node representations in two phases: message passing and updating. These phases are illustrated in Figure 6 (b) and further detailed in Figures 6 (c), (d), and (e). During message passing, nodes receive information from neighboring nodes, expanding their accessible radius. In the updating phase, PaEGNN utilizes the node’s internal messages (composed of $F$ scalars and $F$ vectors) to update its features. To prevent over-smoothingyang2022learning ; yang2022mgraphdta , skip connections are added to each layer.
@@
4.3.1 Periodic cell offset encoding
A notable feature of PaEGNN, distinguishing it from previous modelsschutt2021equivariant ; xie2018crystal , is the explicit differentiation of atoms in various cells to encode PBCs. To achieve this, we define the set $\mathcal{C}=\{-2,-1,0,1,2\}$. We then use this set to generate translated cells with offsets $(k_{1},k_{2},k_{3})\in\mathcal{C}\times\mathcal{C}\times\mathcal{C}$. The translated cells, resulting from the offsets $(k_{1},k_{2},k_{3})\in\mathcal{C}\times\mathcal{C}\times\mathcal{C}$, are generally sufficient to encompass all atoms within a $6\ \text{\AA}$ cutoff distance. We use $(k_{1},k_{2},k_{3})_{ij}$ to denote the cell offset from node $i$ to node $j$, where node $j$ is located in a cell translated by $k_{1}\vec{\bm{l}}_{1}+k_{2}\vec{\bm{l}}_{2}+k_{3}\vec{\bm{l}}_{3}$ relative to node $i$. Let $K_{ij}=(k_{1}+2)+(k_{2}+2)5+(k_{3}+2)25$ be a positive integer that uniquely indexes the cell offset, the sinusoidal positional encodingvaswani2017attention for $K_{ij}$ is computed as:
@@ 4.3.1 Periodic cell offset encoding
The periodic cell offset encoding $\bm{p}(K_{ij})$ explicitly indicates node $j$’s cell location relative to node $i$. This encoding enables the GNN to explicitly recognize the periodic structure, thereby enhancing predictive performance.
@@ 4.3.2 Message passing phase
During this phase, a node $v_{i}$ aggregates messages from neighboring scalars $\bm{x}_{j}$ and vectors $\vec{\bm{x}}_{j}$, forming intermediate scalar and vector variables $\bm{m}_{i}$ and $\vec{\bm{m}}_{i}$ as follows:
@@
$\vec{\bm{m}}_{i}=\sum_{v_{j}\in\mathcal{N}(v_{i})}$ $(\bm{W}_{u}\bm{x}_{j}^{(t)})\circ\gamma_{u}\big{(}\lambda(\|\vec{\bm{r}}_{ji}\|)\oplus\bm{p}(K_{ij})\big{)}\circ\vec{\bm{x}}_{j}^{(t)}$
$+(\bm{W}_{v}\bm{x}_{j}^{(t)})\circ\gamma_{v}\big{(}\lambda(\|\vec{\bm{r}}_{ji}\|)\oplus\bm{p}(K_{ij})\big{)}\circ\frac{\vec{\bm{r}}_{ji}}{\|\vec{\bm{r}}_{ji}\|}$ (5)

### v2

@@ 2.7 DFT validations
Usually, the initial crystal structure may deviate from or be close to the final relaxed structure. To demonstrate the efficacy and robustness of DeepRelax, we perform DFT validations on two types of initial structures: those from the X-Mn-O dataset, which exhibit large deviations from the DFT-relaxed state, and those from the MP dataset, which are generally closer to their DFT-relaxed structures, as illustrated in Suppl. Fig. 1. The detailed settings for the DFT calculations are provided in Subsection 4.9.
In the first experiment, we evaluated our model’s predictive capability under challenging conditions using the X-Mn-O dataset. We filtered out unrelaxed structures from the X-Mn-O test set that are structurally similar to their DFT-relaxed counterparts using Pymatgen’s “Structure_matcher” function. From the remaining test set ($N=1007$), we randomly selected 100 samples. Fig. 4(a) shows the deviation distribution for the selected unrelaxed structures, which closely aligns with that of the complete test set, thus confirming the representativeness of the selected subset. Subsequently, we employed DeepRelax to predict the relaxed structures for these samples. Fig. 4(b) shows box plots of the energy distributions for the unrelaxed, DFT-relaxed, and DeepRelax-predicted structures. The energy distributions of the DeepRelax-predicted and DFT-relaxed structures show similar medians and interquartile ranges, validating the model’s accuracy in predicting energetically favorable structures. The MAE in energy is significantly reduced by an order of magnitude from 32.51 to 5.97.
In the second experiment, we tested whether DeepRelax remains effective with structures starting from a relatively rational initial unrelaxed state using the Materials Project dataset. Here, we again randomly selected 100 samples from the test set. Fig. 4(c) shows the deviation distribution for these samples. The energies of the unrelaxed, DFT-relaxed, and DeepRelax-predicted structures were calculated using DFT. Fig. 4(d) shows that the predicted structures feature an energy distribution nearly identical to that of the DFT-relaxed structures, demonstrating the model’s effectiveness in handling relatively rational initial unrelaxed structures.
Besides the energy indicator, we further demonstrated our model’s effectiveness using the number of residual optimizing steps required for DFT relaxation. Specifically, we randomly selected 20 structures from the test set of the point-defect dataset, with their deviation distribution as shown in Fig. 4(e). We then conducted DFT calculations starting from the unrelaxed and DeepRelax-predicted structures, respectively. As shown in Fig. 4(f), starting DFT relaxation from the DeepRelax-predicted structures significantly reduces the number of required ionic steps, which is also robust.
@@ 2.8 Analysis of uncertainty
A critical challenge in integrating artificial intelligence (AI) into material discovery is establishing trustworthy AI models. Current deep learning models typically offer accurate predictions only within the chemical space covered by their training datasets, known as the applicability domain yu2022uncertainty . Predictions for samples outside this domain can be questionable. Thus, uncertainty quantification has become critical for AI models by quantifying prediction confidence levels, thereby aiding researchers in decision-making and experimental planning.
Figure 5: Uncertainty quantification. Hexagonal binning plots comparing system-level uncertainty with distance MAE ($\rm\AA$) for the (a) X-Mn-O, (b) MP, and (c) C2DB datasets. (d) and (e) illustrate the bond-level uncertainty for each predicted pairwise distance in $\rm Sr_{2}Mn_{2}O_{4}$ and $\rm Mg_{1}Mn_{1}O_{3}$, respectively, demonstrating the correlation between distance prediction errors and their associated bond-level uncertainties. Source data are provided with this paper.
To validate the efficacy of our proposed uncertainty quantification in reflecting the confidence level of model predictions, we compute Spearman’s rank correlation coefficient between the total predicted distance error and its associated system-level uncertainty. Fig. 5(a)-(c) show the hexagonal binning plots of system-level uncertainty against total distance MAE for the X-Mn-O, MP, and C2DB datasets, respectively. Correlation coefficients of 0.95, 0.83, and 0.88 for these datasets demonstrate a strong correlation between predicted error and predicted system-level uncertainty. Fig. 5(d)-(e) present the bond-level uncertainty visualization for two predicted structures, illustrating the correlation between predicted bond length error and associated bond-level uncertainty. These results indicate that the model’s predicted uncertainty is a good indicator of the predicted structure’s accuracy.
Table 4: Ablation study to investigate the impact of unit cell offset encoding (UCOE) and bond-level data uncertainty (BLDU) estimation on model performance. The performances are evaluated by the MAE of coordinates ($\rm\AA$), bond length ($\rm\AA$), lattice ($\rm\AA$), and cell volume (Å$\rm\AA^{3}$) between the predicted and DFT-relaxed structures
@@
DeepRelax (UCOE) 0.121 0.147 0.063 3.563
DeepRelax (BLDU) 0.142 0.171 0.064 3.539
@@
The best performance in each metric is highlighted in bold.
@@ 2.9 Ablation study
DeepRelax’s technical contributions are twofold: it utilizes UCOE for handling PBCs explicitly, and it employs a method for estimating bond-level data uncertainty to encourage the model to capture a more comprehensive representation of the underlying data distribution.
@@
  • Vanilla: Excludes both UCOE and data uncertainty estimation.
  • DeepRelax (UCOE): Integrates UCOE but omits data uncertainty estimation.
  • DeepRelax (BLDU): Implements bond-level data uncertainty estimation but not UCOE.
Table 4 demonstrates that DeepRelax (UCOE) attains a significant performance enhancement over the Vanilla model, suggesting the UCOE contributes greatly to model performance. On the other hand, DeepRelax (BLDU) shows a more modest improvement, which indicates the added value of data uncertainty estimation. Overall, DeepRelax shows a 25.16% improvement in coordinate MAE and a 20.00% advancement in bond length MAE over the Vanilla model. These comparative results underscore the combined effectiveness of UCOE and data uncertainty estimation in our final DeepRelax model.
@@ 3 Discussion
The rapid advancement of generative models like CDVAE xie2021crystal , PGCGM zhao2023physics , and MatterGen zeni2023mattergen , has opened avenues for the prolific generation of hypothetical materials with potentially desirable properties, such as 2.2 million new materials recently discovered by Google DeepMind. Clearly, it is impossible to relax such a huge number of structures using the traditional ab initio method, and it is also very difficult using the iterative ML relax models. For example, we further compare the efficiency between DeepRelax and M3GNet, a leading iterative ML model. DeepRelax offers a substantial speed advantage, being approximately 100 times faster than M3GNet (see Suppl. Table 3). Based on this estimation, to relax the 2.2 million new materials, our DeepRelax model only needs around 100 hours or 4 days, while M3GNet will take around 400 days. Moreover, our DeepRelax model supports parallel GPU processing, which can further significantly reduce computer time. While there are other direct structure-prediction ML methods, such as DOGSS yoon2020differentiable and Cryslator kim2023structure , detailed comparisons with these methods are provided in Suppl. Note 6. Overall, we introduce a fast, scalable, and trustworthy deep generative model, DeepRelax, for direct structural relaxation. Despite its advancements, opportunities for further improvement remain, which we explore in subsequent discussions.
@@ 4 Methods
4.1 Periodicity in crystals
A crystal can be conceptualized as a periodic arrangement of atoms in 3D space xie2021crystal . This periodicity is typically captured by a unit cell, that effectively represents the crystal structure. Such a unit cell, containing $N$ atoms, can be fully characterized by three components:
@@ 4.1 Periodicity in crystals
where $(k_{1},k_{2},k_{3})$ are the unit cell offsets used to replicate the unit cell across the 3D space.
Figure 6: The architecture of PaEGNN. (a) Illustration of the multi-graph representation designed to capture atomic interactions across cell boundaries in periodic structures. (b) Message passing that collects messages from a node’s neighbors and message updating that updates node representations using a node’s internal states. (c) Overview of PaEGNN, comprising four layers, each with message passing and message updating phases, taking unit cell offset integer $K_{ji}$, initial vector $\vec{\bm{x}}_{i}^{(0)}=\vec{\bm{0}}$, initial scalar $\bm{x}_{i}^{(0)}=E(z_{i})$, and relative position $\vec{\bm{r}}_{ij}$ as inputs and outputting the final vector $\vec{\bm{x}}_{i}^{(T)}$ and scalar $\bm{x}_{i}^{(T)}$ representations. (d) During the message passing phase, a node $v_{i}$ aggregates messages from neighboring vectors $\vec{\bm{x}}_{j}^{(t)}$ and scalars $\bm{x}_{j}^{(t)}$, forming intermediate vector and scalar variables $\vec{\bm{m}}_{i}$ and $\bm{m}_{i}$. (e) The message updating phase integrates the $F$ vectors and $F$ scalars within $\vec{\bm{m}}_{i}$ and $\bm{m}_{i}$ to generate updated vector $\vec{\bm{x}}_{i}^{(t+1)}$ and scalar $\bm{x}_{i}^{(t+1)}$.
4.2 Multi-graph representation for crystal structures
Multi-graphs offer an intuitive way to represent crystal structures under periodic boundary conditions (PBCs) xie2021crystal , as depicted in Fig. 6(a). These graphs can be effectively processed by GNNs through graph convolutions or message passing, which simulate many-body interactions li2024jmi ; musaelian2023learning ; pablo2023fast ; li2022deep ; gong2023general ; zhong2023transferable ; zhong2024universal ; xie2018crystal ; park2020developing ; schutt2017schnet ; chen2019graph ; gasteiger_dimenet_2020 ; choudhary2021atomistic ; chen2022universal ; deng2023chgnet ; unke2021spookynet ; batzner20223 ; banik2023cegann ; unke2019physnet ; zhang2022interpretable ; satorras2021n ; omee2022scalable ; schutt2021equivariant ; haghighatlari2022newtonnet ; han2024survey . Formally, we define a multi-graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ to encode these periodic structures. Here, $\mathcal{V}=\{v_{1},...,v_{N}\}$ represents the set of nodes (atoms), and =$\mathcal{E}=\{e_{ij,(k_{1},k_{2},k_{3})}|i,j\in\{1,...,N\},k_{1},k_{2},k_{3}\in\mathbb{Z}\}$ signifies the set of edges (bonds). The edge $e_{ij,(k_{1},k_{2},k_{3})}$ denotes a directed connection from node $v_{i}$ in the original unit cell to node $v_{j}$ in the unit cell translated by $k_{1}\vec{\bm{l}}_{1}+k_{2}\vec{\bm{l}}_{2}+k_{3}\vec{\bm{l}}_{3}$. Nodes are interconnected with their nearest neighbors within a cutoff distance $D$ (set to $6\ \rm\AA$ in our study).
@@ 4.3 Periodicity-aware equivariant graph neural network
PaEGNN iteratively updates node representations in two phases: message passing and updating. These phases are illustrated in Fig. 6(b) and further detailed in Fig. 6(c)-(e). During message passing, nodes receive information from neighboring nodes, expanding their accessible radius. In the updating phase, PaEGNN utilizes the node’s internal messages (composed of $F$ scalars and $F$ vectors) to update its features. To prevent over-smoothing yang2022learning ; yang2022mgraphdta , skip connections are added to each layer.
@@
4.3.1 Unit cell offset encoding
A notable feature of PaEGNN, distinguishing it from previous models schutt2021equivariant ; xie2018crystal , is the explicit differentiation of atoms in various translated unit cells to encode PBCs. To achieve this, we define the set $\mathcal{C}=\{-2,-1,0,1,2\}$. We then use this set to generate translated unit cells with offsets $(k_{1},k_{2},k_{3})\in\mathcal{C}\times\mathcal{C}\times\mathcal{C}$. The translated unit cells, resulting from the offsets $(k_{1},k_{2},k_{3})\in\mathcal{C}\times\mathcal{C}\times\mathcal{C}$, are generally sufficient to encompass all atoms within a $6\ \text{\AA}$ cutoff distance. We use $(k_{1},k_{2},k_{3})_{ij}$ to denote the unit cell offset from node $v_{i}$ to node $v_{j}$, where node $v_{j}$ is located in a unit cell translated by $k_{1}\vec{\bm{l}}_{1}+k_{2}\vec{\bm{l}}_{2}+k_{3}\vec{\bm{l}}_{3}$ relative to node $v_{i}$. Let $K_{ij}=(k_{1}+2)+(k_{2}+2)5+(k_{3}+2)25$ be a positive integer that uniquely indexes the unit cell offset, the sinusoidal positional encoding vaswani2017attention for $K_{ij}$ is computed as:
@@ 4.3.1 Unit cell offset encoding
The unit cell offset encoding $\bm{p}(K_{ij})$ explicitly encodes the relative position of the unit cells in which the two nodes, $v_{i}$ and $v_{j}$, are located. This encoding enables the GNN to explicitly recognize the periodic structure, thereby enhancing predictive performance.
@@ 4.3.2 Message passing phase
During this phase, a node $v_{i}$ aggregates messages from neighboring scalars $\bm{x}_{j}^{(t)}$ and vectors $\vec{\bm{x}}_{j}^{(t)}$, forming intermediate scalar and vector variables $\bm{m}_{i}$ and $\vec{\bm{m}}_{i}$ as follows:
@@
$\vec{\bm{m}}_{i}=\sum_{v_{j}\in\mathcal{N}(v_{i})}$ $(\bm{W}_{u}\bm{x}_{j}^{(t)})\circ\gamma_{u}\big{(}\lambda(\|\vec{\bm{r}}_{ji}\|)\oplus\bm{p}(K_{ji})\big{)}\circ\vec{\bm{x}}_{j}^{(t)}$
$+(\bm{W}_{v}\bm{x}_{j}^{(t)})\circ\gamma_{v}\big{(}\lambda(\|\vec{\bm{r}}_{ji}\|)\oplus\bm{p}(K_{ji})\big{)}\circ\frac{\vec{\bm{r}}_{ji}}{\|\vec{\bm{r}}_{ji}\|}$ (5)